# Time-domain study of coupled collective excitations in quantum materials

### v1

@@ IV.2 Phonon-phonon coupling
The lattice structure often dictates the electronic, optical, and mechanical properties of quantum materials. The collective excitations of atomic vibrations, or phonons, hence play an important role in determining these physical attributes as well. Under equilibrium conditions, phonons and their couplings can be characterized by techniques such as IR spectroscopy 160, Raman spectroscopy 161, and inelastic neutron, X-ray, or electron scatterings 162. Though these static methods are powerful, a lack of precise control over a particular phonon mode impedes us from revealing how individual phonon modes couple to other quasiparticles in solids. In this regard, ultrafast techniques allow us to tackle this challenge by extending the measurement into the time domain. For example, advances in generating intense mid-IR and THz pulses with either broad band or tunable narrow-band frequencies bring the paradigm of nonlinear phononics 163, 47. By selectively exciting particular phonon modes using tailored low-photon-energy pulses, we can study phonon-phonon couplings and engineer crystal structure in a nonequilibrium state.
@@
The first experimental demonstration of nonlinear phononics was reported in 2011 by Först and coworkers 164. After the IR mode in a bulk single crystal of La0.7Sr0.3MnO3 was excited by 14.3-m light pulses, the reflectivity exhibited an oscillation at 1.2 THz, as shown by the dark curve in Fig. 5a. The oscillation was attributed to the Raman mode that is associated with oxygen octahedral rotation in this perovskite. Moreover, the amplitude of the mode showed a strong pump wavelength dependence and reached a maximum when it was resonant with the mode (see Fig. 5b). This intriguing phenomenon can be described by an anharmonic coupling between the IR-active mode and the Raman-active mode, as illustrated in Fig. 4. During this process, a mid-IR light pulse resonantly couples to the dipole of the IR-active phonon and sets off coherent oscillations of atoms around their equilibrium positions (Fig. 4a). The anharmonic coupling term () applies a directional force (Fig. 4b) on the Raman mode (Fig. 4c), thus driving its oscillations. For a more detailed description of this mechanism such as symmetry considerations during anharmonic interactions, we refer readers to refs. 165, 47. In contrast to the resonant excitation condition, 5.8-THz oscillations were observed when La0.7Sr0.3MnO3 was photoexcited by a 1.5-m pulse (Fig. 5a, light curve). This phonon mode was attributed to the mode, which was induced by the more conventional mechanism involving a displacive excitation of coherent phonons 15.
@@
This work by Först et al. inspired ideas to engineer crystals using resonant mid-IR or terahertz excitations. This nonlinear phononic mechanism has since been utilized to manipulate properties in many perovskite systems, such as controlling an anisotropic strain wave in LaAlO3 166, as well as changing ferroelectric polarizations in systems such as LiNbO3 (refs. 167, 168), SrTiO3 (refs. 169, 170, 171), and BiFeO3 (ref. 172). Besides the effect on the lattice and hence the coupled polar orders, nonlinear phononics have also been shown to influence electronic and magnetic properties 47. A particularly intriguing case concerns the apparent resemblance to a transient superconducting state when resonant excitation of IR-active phonons of specific wavelengths are driven to a large amplitude 173, 174, 175, 176, 177, 178. However, other works have suggested that certain superconducting-like behavior may be a manifestation of the excitation of quasiparticles with a low scattering rate 179, 180, and that care needs to be taken in processing the transient optical data for drawing precise conclusions about the nonequilibrium state 181, 182, 183. This potential of light-induced superconductivity remains an intensely studied area with many interesting, open questions to be answered.
@@
Compared with ultrafast optical spectroscopy, time-resolved X-ray diffraction can directly quantify the changes in a crystal structure, thus offering a powerful alternative to study phonon-phonon couplings 36, 35. Here, we highlight several experimental works on SrTiO3, a quantum paraelectric material where the competition between ferroelectric and antiferrodistortive instabilities prevents it from entering a ferroelectric phase at low temperature. Using THz pump pulses, the direct control of its ferroelectric soft phonon mode has been intensely studied and several nonlinear phononic effects have been observed 169, 170, 171. For example, phonon up-conversion in SrTiO3 was demonstrated in a THz pump, X-ray diffraction probe experiment 163. Kozina et al. used an intense THz pump pulse with a spectral range of 0.2–2.5 THz (red region in Fig. 5c), which overlapped with the IR-active TO1 soft mode in SrTiO3 at 100 K (the yellow peak). Interestingly, the oscillatory Bragg peak intensity observed in X-ray diffraction not only showed the frequency of the driving THz field but also exhibited several high frequencies that correspond to the TO2 and TO3 modes (turquoise and purple peaks in Fig. 5c). Upon a temperature increase where the hardened TO1 mode falls outside the THz excitation spectrum, no signatures of the TO2 and TO3 modes are seen, providing compelling evidence for phonon frequency up-conversion from the soft TO1 mode to these high-frequency modes.
@@
Beyond the proof-of-principle demonstration by Kozina et al., nonlinear phononic effects have also been used to induce ferroelectricity in unstrained SrTiO3 under ambient pressure. In 2019, Li and coworkers reported that SrTiO3 can be driven to a transient ferroelectric phase via intense THz pumping as evidenced by the second harmonic signal of a 800-nm probe pulse 169, although later experiments performed on a cousin compound KTaO3 suggested that the THz-induced transient second-harmonic signal may come from defect-induced local polar structures without any long-range ferroelectric ordering 185. A contemporary work in 2019 reported a metastable ferroelectric phase of SrTiO3, which was induced by a mid-IR pump pulse instead of a THz pulse 170. In this experiment, Nova and coworkers resonantly drove the high-frequency IR-active mode in SrTiO3 at 4 K, and detected a metastable noncentrosymmetric phase via second-harmonic spectroscopy. The experiment was interpreted via anharmonic phonon-phonon interactions between the pulse-induced IR-active mode and acoustic strains, the latter of which suppressed the antiferrodistortive distortion and hence stabilized the ferroelectric phase. This mechanism was further clarified by a subsequent work using time-resolved X-ray diffusing scattering 171. The antiferrodistortive fluctuations can lead to a significant diffuse scattering intensity at the point in the Brillouin zone, as shown in the equilibrium X-ray detector image in Fig. 5d. At 135 K, upon photoexcitation with a 17-THz pulse that resonantly drives the IR-active mode, X-ray scattering intensity at shows a sudden increase followed by a decay (Fig. 5e). The sudden intensity increase is a result of the enhancement of antiferrodistortive fluctuations induced by their anharmonic coupling with the resonantly excited high-frequency IR-active mode. The same IR-active mode can also couple to low-frequency acoustic modes, which induce a strain that dominates at long time delays ( ps) and suppresses the antiferrodistortive fluctuations, resulting in the decrease of X-ray intensity at the point. Though the X-ray experiment was performed at 135 K that is above the antiferrodistortive transition, the authors expected the same mechanism would remain valid to explain the metastable ferroelectric property of SrTiO3 at 4 K, where a suppressed antiferro-distortion at long times leads to the emergence of a ferroelectric phase.
@@
Aside from the standard pump-probe scheme employed in transient optical spectroscopy and X-ray scattering, 2D THz spectroscopy offers another route to determine the interaction pathway between different quasiparticles, showing particular relevance in revealing the phonon-phonon coupling process 29. In typical 2D THz spectroscopy setups, two intense THz pump pulses are used to induce nonlinear responses with a time delay between the two pulses, as shown in Fig. 5f. A third optical or near-infrared probe pulse with time delay is then applied to record the optical response of the sample. By performing a 2D Fourier transform on the time-domain data with varying and , spectra as a function of THz excitation frequency and detection frequency can be generated, yielding insights into the interaction process between different modes. As a recent example, 2D THz spectroscopy was employed to reveal the energy-flow pathway in a topological antiferromagnet MnBi2Te4 (ref. 184). The normalized 2D Fourier transform spectrum of the nonlinear signal is shown in Fig. 5g, while Fig. 5h shows three simulated spectra based on the different coupling pathways of exciting the Raman-active phonon: two-photon absorption (“photonic”), nonlinear phononics (“phononic”), and a combination of the two (“photophononic”) . Careful comparison with the experimental observation in Fig. 5g pinpoints a photophononic pathway as the most probable scenario, where the THz electric field first excites the IR-active mode and then interacts with this mode to further excite the Raman-active mode, as summarized in Fig. 5i.
@@ IV.3 Phonon-magnon coupling
Novel properties in quantum materials often arise from the coupling between the spin and lattice degrees of freedom, for example, the Einstein-de Haas effect 190, the Barnett effect 191, 192, magnetostriction 193, and piezomagnetism 194, 195. In certain contexts, their strong coupling implies a unique spin configuration dictated by the crystalline lattice, and one prototypical example is the realization of quantum spin liquid predicted in a frustrated geometry such as in a triangular or kagome lattice 196.
@@
While effects due to strong spin-lattice coupling have been extensively investigated in thermal equilibrium 197, ultrafast techniques allow us to probe the exact energy-flow pathway and unveils how angular momentum is transferred between the electron spin and the crystalline lattice. For instance, ferromagnetic systems are known to lose their net magnetization well within a picosecond when they are subjected to an intense ultrafast laser pulse 198, but how the conservation of total angular momentum is maintained at this timescale has been a long-standing puzzle. Recent works using time-resolved X-ray and electron diffraction offered some clues to this question 199, 200, where circularly-polarized high-frequency phonons are first excited to absorb the angular momentum, which are then transferred to low-frequency shear acoustic modes and a macroscopic rotation of the sample. Interestingly, acoustic shearing is similarly reported during ultrafast demagnetization of a van der Waals antiferromagnet whose net spin angular momentum is zero 201, 21, 202, and a detailed description of the angular momentum flow in the antiferromagnetic case remains an open question.
@@
Strong spin-lattice couplings can manifest in the form of magnon-phonon coupling, where a hybridized magnon-phonon mode, magnon polarons, can readily be detected in equilibrium. Similar to phonon-polaritons discussed in Sec. IV.1, one characteristic feature of magnon polarons is the avoided crossing of energy levels of the constituent magnon and phonon modes. An example of avoided crossings is shown in Fig. 6a for a layered antiferromagnet FePSe3, measured by magneto-infrared transmission spectroscopy 186. In the detected spectral window of FePSe3, while the constituent phonons are individually chiral with opposite chiralities, their near degeneracy leads to no net chirality. However, the coupling to chiral magnons lifts such degeneracy, resulting in chiral magnon polarons that were demonstrated by magento-Raman spectroscopy. The hybridization of magnon and phonon can also be directly measured in the time domain using ultrafast techniques. For instance, in a cousin van der Waals antiferromagnet FePS3, two Raman-active optical phonons at 3.2 and 4.8 THz can be excited by a near-infrared pump pulse, evidenced by the coherent oscillations in the polarization rotation signal of the probe pulse (Fig. 6b) 187. As the applied magnetic field increases, the 3.2 THz phonon splits into two phonon-magnon modes due to phonon-magnon coupling (Fig. 6c), which is reminiscent of the hybridization-induced peak splittings detected in equilibrium 203, 204, 205.
@@
In the study of phonon-magnon coupling, what sets nonequilibrium experiments apart from their equilibrium counterpart is the coherence of the dynamical motion of both atoms and spins, leading to a number of interesting phenomena. For example, circularly polarized phonons can be coherently excited, leading to an effective magnetic field that can in turn excite coherent spin precession. In an early demonstration of this effect, Nova and coworkers used a 20-THz mid-infrared pump to resonantly drive the in-plane and phonons in an antiferromagnetic insulator ErFeO3 (ref. 188), whose phonon eigenvectors are sketched in Fig. 7a. Due to the orthorhombic distortion in ErFeO3, these two phonons exhibit different eigenfrequencies and oscillator strengths. When the modes are excited simultaneously, the two modes begin with a nonzero relative phase and evolve at different rates, resulting in an elliptical motion of oxygen atoms, as illustrated in the bottom panel of Fig. 7a. The elliptically polarized phononic field mimics the application of a magnetic field and results in a spin precession, leading to the excitation of a coherent magnon at 0.75 THz frequency, which is readily detected in the Fourier spectrum of the pump-induced Faraday rotation signal (Fig. 7b). The elliptically polarized phonons are otherwise known as chiral phonons, which have attracted both theoretical 207, 208 and experimental 209, 76, 70, 69 interests because of their ability to create an effective magnetic field that paves the way for “chiral phonomagnetism” in magnetization-based devices.
@@
Coherent magnons can also be excited by phonon-induced lattice distortions that modify the exchange interaction between neighboring atoms, a mechanism recently demonstrated in a closely related antiferromagnet DyFeO3 by Afanasiev and colleagues 189. In this material, the highest frequency IR-active transverse optical phonon mode () is associated with a periodic stretching of the Fe-O bond. Using a tailored terahertz pump pulse, this phonon can be resonantly driven to a large amplitude and couple to a finite lattice distortion along the coordinate of a Raman-active mode via the nonlinear phononics pathway discussed in Sec. IV.2 (see Fig. 4). This lattice distortion involves an antipolar motion of the Dy3+ ions, resulting in a rectification of the Fe-Dy exchange integral that can reach approximately 1–2 eV per unit cell according to density functional theory calculations, thus setting off a coherent magnon oscillation. Depending on whether DyFeO3 is in the antiferromagnetic state ( K) or the weakly ferromagnetic state ( K), distinct magnon modes are induced by the resonant driving of the same optical phonon (Fig. 7c,d). Importantly, the magnon mode frequencies under the resonant driving condition (orange peak in the insets) are clearly distinct from those under off-resonant pumping (black peak in the insets), the latter of which are excited via impulsive stimulated Raman scattering 210. Leveraging the modification of exchange interaction under a resonant pumping condition, Afanasiev and colleagues further showed that with a very large fluence of the resonant pump, an ultrafast magnetic transition from an antiferromagnetic state to a weakly ferromagnetic state can be induced coherently, a transition that proceeded much faster compared to an above-bandgap excitation scheme.
@@
While the preceding examples illustrated how coherent magnons can be induced by resonant excitation of coherent phonons, the reverse pathway from magnons to phonons has also been demonstrated. In an experiment conducted on antiferromagnetic CoF2, Mashkovich et al. used an intense, nearly single-cycle terahertz pulse to pump the crystal below its Néel temperature, where the magnetic field of the pulse resonantly excited a magnon mode (Fig. 8a) 206. After approximately 10 ps, the amplitude of the magnon decreases, followed by a persistent phonon oscillation over the next tens of picoseconds. By contrast, the magnon and phonon oscillations both disappeared above the Néel temperature (Fig. 8b), strongly suggesting that the phonons are induced by the magnons and not vice versa. The researchers further performed 2D terahertz spectroscopy to demonstrate the causal relation that the nonlinear excitation of phonons was mediated by a coherent magnon, as shown by the Fourier spectrum of the nonlinear amplitude in Fig. 8c. Specifically, a peak was observed only when the excitation frequency coincided with the magnon mode and the detection frequency coincided with the phonon mode; on the other hand, no peak was seen above the noise level when one swapped the excitation and detection frequency. The magnon-mediated excitation pathway for the coherent phonon is summarized in Fig. 8e. In the first step, an intense terahertz field resonantly populates the coherent magnonic state at frequency , creating an intermediate state. In the second step, another terahertz photon at frequency interacts with this intermediate state and coherently excites the phonon.

### v2

@@ IV.2 Phonon-phonon coupling
The lattice structure often dictates the electronic, optical, and mechanical properties of quantum materials. The collective excitations of atomic vibrations, or phonons, hence play an important role in determining these physical attributes as well. Under equilibrium conditions, phonons and their couplings can be characterized by techniques such as IR spectroscopy 164, Raman spectroscopy 165, and inelastic neutron, X-ray, or electron scatterings 166. Though these static methods are powerful, a lack of precise control over particular phonon modes impedes us from revealing how individual phonon modes couple to other quasiparticles in solids. In this regard, ultrafast techniques allow us to tackle this challenge by extending the measurement into the time domain. For example, advances in generating intense mid-IR and THz pulses with either broadband or tunable narrow-band frequencies have been instrumental in the emergence of the field of nonlinear phononics 167, 47. By selectively exciting particular phonon modes using tailored low-photon-energy pulses, we can study phonon-phonon couplings and engineer crystal structure in a nonequilibrium state.
@@
The first experimental demonstration of nonlinear phononics was reported in 2011 by Först and coworkers 167. After the IR mode in a bulk single crystal of La0.7Sr0.3MnO3 was excited by 14.3-m light pulses, the reflectivity exhibited oscillations at 1.2 THz, as shown by the dark curve in Fig. 5a. The oscillations were attributed to the Raman mode that is associated with oxygen octahedral rotation in this perovskite. Moreover, the amplitude of the mode showed a strong pump wavelength dependence and reached a maximum when it was resonant with the mode (see Fig. 5b). This intriguing phenomenon can be described by an anharmonic coupling between the IR-active mode and the Raman-active mode, as illustrated in Fig. 4. During this process, a mid-IR light pulse resonantly couples to the dipole of the IR-active phonon and sets off coherent oscillations of atoms around their equilibrium positions (Fig. 4a). The anharmonic coupling term () applies a directional force (Fig. 4b) on the Raman mode (Fig. 4c), thus driving its oscillations. For a more detailed description of this mechanism such as symmetry considerations during anharmonic interactions, we refer readers to refs. 168, 47. In contrast to the resonant excitation condition, 5.8-THz oscillations were observed when La0.7Sr0.3MnO3 was photoexcited by a 1.5-m pulse (Fig. 5a, light curve). This phonon mode was attributed to the mode, which was induced by the more conventional mechanism involving a displacive excitation of coherent phonons 15.
@@
This work by Först et al. inspired ideas to engineer crystals using resonant mid-IR or terahertz excitations. This nonlinear phononic mechanism has since been utilized to manipulate properties in many perovskite systems, such as controlling an anisotropic strain wave in LaAlO3 (ref. 169), as well as changing ferroelectric polarizations in systems such as LiNbO3 (refs. 170, 171), SrTiO3 (refs. 172, 173, 174), and BiFeO3 (ref. 175). Besides the effect on the lattice and hence coupled polar orders, nonlinear phononics has also been shown to influence electronic and magnetic properties 47. A particularly intriguing case concerns the apparent resemblance to a transient superconducting state when IR-active phonons of specific wavelengths are driven to a large amplitude by resonant excitation 176, 177, 178, 179, 180, 181. However, other works have suggested that certain superconducting-like behavior may be a manifestation of the excitation of quasiparticles with a low scattering rate 182, 183. Therefore, care needs to be taken in processing the transient optical data for drawing precise conclusions about the nonequilibrium state 184, 185, 186. This potential of light-induced superconductivity remains an intensely studied area with many interesting, open questions to be answered.
@@
Compared with ultrafast optical spectroscopy, time-resolved X-ray diffraction can directly quantify the changes in a crystal structure, thus offering a powerful alternative to study phonon-phonon couplings 36, 35. Here, we highlight several experimental works on SrTiO3, a quantum paraelectric material where the competition between ferroelectric and antiferrodistortive instabilities prevents it from entering a ferroelectric phase at low temperature. Using THz pump pulses, the direct control of its ferroelectric soft phonon mode has been intensely studied and several nonlinear phononic effects have been observed 172, 173, 174. For example, phonon up-conversion in SrTiO3 was demonstrated in a THz pump, X-ray diffraction probe experiment 187. Kozina et al. used an intense THz pump pulse with a spectral range of 0.2–2.5 THz (red region in Fig. 5c), which overlapped with the IR-active TO1 soft mode in SrTiO3 at 100 K (the yellow peak). Interestingly, the oscillatory Bragg peak intensity observed in X-ray diffraction not only showed the frequency of the driving THz field but also exhibited several high frequencies that correspond to the TO2 and TO3 modes (turquoise and purple peaks in Fig. 5c). When temperature increased such that the hardened TO1 mode fell outside the THz excitation spectrum, no signatures of the TO2 and TO3 modes were seen, providing compelling evidence for phonon frequency up-conversion from the soft TO1 mode to these high-frequency modes.
@@
Beyond the proof-of-principle demonstration by Kozina et al., nonlinear phononic effects have also been used to induce ferroelectricity in unstrained SrTiO3 under ambient pressure. In 2019, Li and coworkers reported that SrTiO3 can be driven to a transient ferroelectric phase via intense THz pumping as evidenced by the second harmonic signal of a 800-nm probe pulse 172, although later experiments performed on a cousin compound KTaO3 suggested that the THz-induced transient second-harmonic signal may come from defect-induced local polar structures without any long-range ferroelectric ordering 189. A contemporary work in 2019 reported a metastable ferroelectric phase of SrTiO3, which was induced by a mid-IR pump pulse instead of a THz pulse 173. In this experiment, Nova and coworkers resonantly drove the high-frequency IR-active mode in SrTiO3 at 4 K, and detected a metastable noncentrosymmetric phase via second-harmonic spectroscopy. The results were interpreted via anharmonic phonon-phonon interactions between the pulse-induced IR-active mode and acoustic strain, the latter of which suppressed the antiferrodistortive distortion and hence stabilized the ferroelectric phase. This mechanism was further clarified by a subsequent work using time-resolved X-ray diffusing scattering 174. The antiferrodistortive fluctuations can lead to a significant diffuse scattering intensity at the point in the Brillouin zone, as shown in the equilibrium X-ray detector image in Fig. 5d. At 135 K, upon photoexcitation with a 17-THz pulse that resonantly drives the IR-active mode, the X-ray scattering intensity at shows a sudden increase followed by a decay (Fig. 5e). The sudden intensity increase is a result of the enhancement of antiferrodistortive fluctuations induced by their anharmonic coupling with the resonantly excited high-frequency IR-active mode. The same IR-active mode can also couple to low-frequency acoustic modes, which induce a strain that dominates at long time delays ( ps) and suppresses the antiferrodistortive fluctuations, resulting in the decrease of X-ray intensity at the point. Though the X-ray experiment was performed at 135 K, which is above the antiferrodistortive transition, the authors expect the same mechanism would remain valid to explain the metastable ferroelectric property of SrTiO3 at 4 K, where a suppressed antiferro-distortion at long times leads to the emergence of a ferroelectric phase.
@@
Aside from the standard pump-probe scheme employed in transient optical spectroscopy and X-ray scattering, 2D THz spectroscopy offers another route to determine the interaction pathway between different quasiparticles, showing particular relevance in revealing phonon-phonon coupling 29. In typical 2D THz spectroscopy setups, two intense THz pump pulses are used to induce nonlinear responses with a time delay between the two pulses, as shown in Fig. 5f. A third optical or near-infrared probe pulse with time delay is then applied to record the optical response of the sample. By performing a 2D Fourier transform on the time-domain data with varying and , spectra as a function of THz excitation frequency and detection frequency can be generated, yielding insights into the interaction process between different modes. As a recent example, 2D THz spectroscopy was employed to reveal the energy-flow pathway in a topological antiferromagnet MnBi2Te4 (ref. 188). The normalized 2D Fourier transform spectrum of the nonlinear signal is shown in Fig. 5g, while Fig. 5h shows three simulated spectra based on the different coupling pathways of exciting the Raman-active phonon: two-photon absorption (“photonic”), nonlinear phononics (“phononic”), and a combination of the two (“photophononic”) . Careful comparison with the experimental observation in Fig. 5g pinpoints a photophononic pathway as the most probable scenario, where the THz electric field first excites the IR-active mode and then interacts with this mode to further excite the Raman-active mode, as summarized in Fig. 5i.
@@ IV.3 Phonon-magnon coupling
Novel properties in quantum materials often arise from the coupling between the spin and lattice degrees of freedom, for example, the Einstein-de Haas effect 194, the Barnett effect 195, 196, magnetostriction 197, and piezomagnetism 198, 199. In certain contexts, their strong coupling implies a unique spin configuration dictated by the crystalline lattice, and one prototypical example is the realization of quantum spin liquid predicted in a frustrated geometry such as in a triangular or kagome lattice 200.
@@
While effects due to strong spin-lattice coupling have been extensively investigated in thermal equilibrium 201, ultrafast techniques allow us to probe the exact energy-flow pathway and reveal how angular momentum is transferred between the electron spin and the crystalline lattice. For instance, ferromagnetic systems are known to lose their net magnetization well within a picosecond when they are subjected to an intense ultrafast laser pulse 202, but how the conservation of total angular momentum is maintained at this timescale has been a long-standing puzzle. Recent works using time-resolved X-ray and electron diffraction offered some clues to this question 203, 204, where circularly-polarized high-frequency phonons were first excited to absorb the angular momentum, which was then transferred to low-frequency shear acoustic modes and a macroscopic rotation of the sample. Interestingly, acoustic shearing has been similarly reported during ultrafast demagnetization of a van der Waals antiferromagnet whose net spin angular momentum is zero 205, 21, 206, and a detailed description of the angular momentum flow in the antiferromagnetic case remains an open question.
@@
Strong spin-lattice couplings can sometimes manifest in the form of magnon-phonon coupling, where hybridized magnon-phonon modes, magnon polarons, can readily be detected in equilibrium. Similar to phonon-polaritons discussed in Sec. IV.1, one characteristic feature of magnon polarons is the avoided crossing of energy levels of the constituent magnon and phonon modes. An example of avoided crossings is shown in Fig. 6a for a layered antiferromagnet FePSe3, measured by magneto-infrared transmission spectroscopy 190. In the detected spectral window of FePSe3, while the constituent phonons are individually chiral with opposite chiralities, their near degeneracy leads to no net chirality. However, the coupling to chiral magnons lifts such degeneracy, resulting in chiral magnon polarons that were demonstrated by magneto-Raman spectroscopy. The hybridization of magnon and phonon can also be directly measured in the time domain using ultrafast techniques. For instance, in a cousin van der Waals antiferromagnet FePS3, two Raman-active optical phonons at 3.2 and 4.8 THz can be excited by a near-infrared pump pulse, evidenced by the coherent oscillations in the polarization rotation signal of the probe pulse (Fig. 6b) 191. As the applied magnetic field increases, the 3.2 THz phonon splits into two phonon-magnon modes due to phonon-magnon coupling (Fig. 6c), which is reminiscent of the hybridization-induced peak splittings detected in equilibrium 207, 208, 209.
@@
In the study of phonon-magnon coupling, what sets nonequilibrium experiments apart from their equilibrium counterpart is the coherence of the dynamical motion of both atoms and spins, leading to a number of interesting phenomena. For example, circularly polarized phonons can be coherently excited, leading to an effective magnetic field that can in turn excite coherent spin precession. In an early demonstration of this effect, Nova and coworkers used a 20-THz mid-infrared pump to resonantly drive the in-plane and phonons in an antiferromagnetic insulator ErFeO3 (ref. 192), whose phonon eigenvectors are sketched in Fig. 7a. Due to the orthorhombic distortion in ErFeO3, these two phonons exhibit different eigenfrequencies and oscillator strengths. When the modes are excited simultaneously, the two modes begin with a nonzero relative phase and evolve at different rates, resulting in an elliptical motion of oxygen atoms, as illustrated in the bottom panel of Fig. 7a. The elliptically polarized phononic field mimics the application of a magnetic field and results in a spin precession, leading to the excitation of a coherent magnon at 0.75 THz frequency, which is readily detected in the Fourier spectrum of the pump-induced Faraday rotation signal (Fig. 7b). The elliptically polarized phonons are otherwise known as chiral phonons, which have attracted both theoretical 211, 212 and experimental 213, 76, 70, 69 interests because of their ability to create an effective magnetic field that paves the way for “chiral phonomagnetism” in magnetization-based devices.
@@
Coherent magnons can also be excited by phonon-induced lattice distortions that modify the exchange interaction between neighboring atoms, a mechanism recently demonstrated in a closely related antiferromagnet DyFeO3 by Afanasiev and colleagues 193. In this material, the highest-frequency IR-active transverse optical phonon mode () is associated with a periodic stretching of the Fe-O bond. Using a tailored terahertz pump pulse, this phonon can be resonantly driven to a large amplitude and couple to a finite lattice distortion along the coordinate of a Raman-active mode via the nonlinear phononics pathway discussed in Sec. IV.2 (see Fig. 4). This lattice distortion involves an antipolar motion of the Dy3+ ions, resulting in a rectification of the Fe-Dy exchange integral that can reach approximately 1–2 eV per unit cell according to DFT calculations, thus setting off coherent magnon oscillations. Depending on whether DyFeO3 is in the antiferromagnetic state ( K) or the weakly ferromagnetic state ( K), distinct magnon modes are induced by the resonant driving of the same optical phonon (Fig. 7c,d). Importantly, the magnon mode frequencies under the resonant driving condition (orange peak in the insets) are clearly distinct from those under off-resonant pumping (black peak in the insets), the latter of which are excited via impulsive stimulated Raman scattering 214. Leveraging the modification of exchange interaction under a resonant pumping condition, Afanasiev and colleagues further showed that with a very large fluence of the resonant pump, an ultrafast magnetic transition from an antiferromagnetic state to a weakly ferromagnetic state can be induced coherently, a transition that proceeded much faster compared to an above-bandgap excitation scheme.
@@
While the preceding examples illustrated how coherent magnons can be induced by resonant excitation of coherent phonons, the reverse pathway from magnons to phonons has also been demonstrated. In an experiment conducted on antiferromagnetic CoF2, Mashkovich et al. used an intense, nearly single-cycle terahertz pulse to pump the crystal below its Néel temperature, where the magnetic field of the pulse resonantly excited a magnon mode (Fig. 8a) 210. After approximately 10 ps, the amplitude of the magnon decreased, followed by a persistent phonon oscillation over the next tens of picoseconds. By contrast, the magnon and phonon oscillations both disappeared above the Néel temperature (Fig. 8b), strongly suggesting that the phonons are induced by the magnons and not vice versa. The researchers further performed 2D terahertz spectroscopy to demonstrate the causal relation that a coherent magnon mediated the nonlinear excitation of phonons, as shown by the Fourier spectrum of the nonlinear amplitude in Fig. 8c. Specifically, a peak was observed only when the excitation frequency coincided with the magnon mode and the detection frequency coincided with the phonon mode; on the other hand, no peak was seen above the noise level when one swapped the excitation and detection frequency. The magnon-mediated excitation pathway for the coherent phonon is summarized in Fig. 8d. In the first step, an intense terahertz field resonantly populates the coherent magnonic state at frequency , creating an intermediate state. In the second step, another terahertz photon at frequency interacts with this intermediate state and coherently excites the phonon.